OpenStudy (anonymous):

Derman says I'm halfway through my book if i read another 83 pages, i'll be two thirds done with my book. How many pages are Dermans book? Show work please.

OpenStudy (anonymous):

let consider total pages be x by question, he has read half of it which means he studied x/2 part of book again, as question if he study 83 more pages, he would have finished 2/3 part of book. so, x/2+83= 2/3 of x or, x+166/2 = 2x/3 or, 3x+498 = 4x or, 498 = x

OpenStudy (anonymous):

Thank you I get it now :)

OpenStudy (anonymous):

no mention dude